OpenStudy (anonymous):

Find the coordinates of the foci for the ellipse: (x + 3)^2 /9 + (y - 1)^2 /25 = 1 a. (3, -1) and (3, 5) b. (-3, 3) and (-3, 1) c. (-3, 5) and (-3, -3) d. (1,1) and (-7,1)

OpenStudy (mertsj):

c^2=a^2-b^2 c^2=25-9 c^2 = 16 c= 4,-4 Note that the major axis of this ellipse is vertical since the big number is under y^2. Also note that the center of this ellipse is (-3,1) and c = 4 so 4 units above (-3,1) is (-3,5). That is one of the focal points. The other one is 4 units below (-3,1) which is (-3,-3). So (-3,5) and (-3,-3) or c